OpenStudy (anonymous):

I am given these 2 t/f questions: If you could please help explain them I would appreciate it! a.) Let R denote the set of real numbers. For real numbers a and b, a does not equal 0, the assignments x --> ax+b is a permutation R --> R. Do these permutations form a group under composition of functions? b.)Let H be the set consisting of the following permutations of the plane: All translations and all reflections in any line of the plane. Is H a group under composition of mappings?

OpenStudy (anonymous):

So for something to be a group it needs a set and an operation.

OpenStudy (anonymous):

In this case the set is \(\mathbb R\). What would the operation be?

OpenStudy (anonymous):

composition?

OpenStudy (anonymous):

I wil give a medal please someone help

OpenStudy (anonymous):

anyone? Ill pay $5.00 if I have to.. I have a quiz coming up that is similar to this and would like to understand whats going on..

OpenStudy (anonymous):

@phi @tkhunny could any of you guys help me?

OpenStudy (tkhunny):

Does the suggested operation fit all the requirements of a group? I'd start with Closure.

OpenStudy (anonymous):

yes its closed: let h(x) = ax+b g(x)=cx+d, So g(h(x)) = g(cx+d)=a(cx+d)+b=(ac)x+cb+d ?

OpenStudy (tkhunny):

Well, that's a little funny. Seems like a negative is sneaking in there. Also, you should be MUCH more formal about parameters you introduce. For Real Numbers e, f, c, d, with neither e, nor c equal to zero (0). h(x) = ex+f g(x) = cx+d g(h(x)) = c(h(x)) + d = c(ex+f) + d = ecx + fc + d = (ec)x + (fc+d) h(g(x)) will be identical. If we agree that Real Numbers have Closure, then we are done, since: a = ec and b = fc + d Okay, we're done with Closure The official list is: 1) Closure 2) Associativity 3) Existence of an Identity Element under the Operation 4) Existence of an Inverse for each element. Prove all four? Group!!! Yea! (Okay, it can be a little tedious.)

OpenStudy (anonymous):

so for Associativity: f . g . h= f . (g . h) where . is composition?

OpenStudy (tkhunny):

Not quite. (fg)h = f(gh) which may have been what you meant. It's just more clear.